OpenStudy (anonymous):

Write y = –4x2 – 64x – 265 in vertex form.

OpenStudy (solomonzelman):

I will do it and you tell me what shape it is, okay ?

OpenStudy (anonymous):

Can you do a vertex form in a circle, @SolomonZelman I've been wondering.

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

This won't be a circle, Jesstho

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y=-4x^2-64x-265 }\) \(\normalsize\color{blue}{ y=-4x^2-16x+256-256-265 }\) \(\normalsize\color{blue}{ y=-4(x^2-16x+64)-256-265 }\) \(\normalsize\color{blue}{ y=-4(x^2-16x+64)-521 }\) \(\normalsize\color{red}{ y=-4(x-8)^2-521 }\)

OpenStudy (solomonzelman):

What is it, what do you think ?

OpenStudy (solomonzelman):

What is it, what do you think ?

OpenStudy (anonymous):

I would say a square.

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

\(\normalsize\color{green}{ y=a( \rm {{x}} - h) ^2+k }\) dowsn't remind you anything ?

OpenStudy (anonymous):

wait wait, i know thats something about parbolas

OpenStudy (solomonzelman):

Yes, this is opening down parabola.

OpenStudy (anonymous):

sorry im exhausted ive been cramming the past week and im fried xD

OpenStudy (anonymous):

i dont think its -521 tho

OpenStudy (anonymous):

@SolomonZelman